40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 3 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  16767   Fri Apr 8 16:03:58 2022 ranaUpdateElectronicsRFSoC 2x2 board -- setup for remote work & BALUN saga

Seems like it should be possible to just remove the transformer (aka as a BALUN ... BALanced, UNbalanced), or replace it with a lower frequency part. Its just a usual mini-circuits part. Maybe you can ask Chris Stoughton about this and ask Tommy to checkout some of the RFSoC user forums for how to go to DC.

Quote:
 

After fiddling around with the tone-generators and spectrum analyzer tools in loopback configuration (DAC --> ADC direct connection), we noticed that lower frequency (~ 1 MHz) signals were hardly making it out/back into the board... so we looked at some of the schematics found here and saw that both RF data converters (ADC & DAC) interfaces are AC coupled through a BALUN network in the 10 - 8000 MHz band (see Attachment #1). This is in principle not great news if we want to get this board ready for audio-band DSP.

We decided that while Tommy works on measuring TFs for SHP-200 all the way up to ~ 2 GHz (which is possible with the board as is) I will design and put together an analog modulation/demodulation frontend so we can upconvert all our "slow" signals < 1MHz for fast, wideband DSP. and demodulate them back into the audio band. The BALUN network is pictured in Attachment #2 on the board, I'm afraid it's not very simple to bypass without damaging the PCB or causing some other unwanted effect on the high-speed DSP.

 

  16766   Thu Apr 7 21:15:04 2022 YehonathanUpdateBHDInitial BHD modeling: AS - LO mode matching

Ok, it turns out these optics were purchased on purpose, as this elog shows. Jon considered building a mode-matching telescope with stock optics as an initial step before purchasing the custom optics (referred to as "design" optics in my elog).

I dialed in the new distances between the optics into the .kat file as described in this elog and pushed the changes to the repo. With the new distances, I got mode-matching of 87% for the full IFO and 89% for FPMI so there's probably no need to worry as the mode-matching with these optics was already designed.

Quote:

I begin modeling the initial BHD setup using Finesse. I started with copying C1_w_BHD.kat from the 40m/bhd repo and making changes to reflect the current BHD setup:

1. OMCs were removed.

2. Only 1 PD per BHD port was left.

3. Transmission of PR2 was changed to 2.2%. The PRG was calculated to be ~15.5.

4. Actual RoCs of new optics were dialed in (Yesterday me and Paco went into the cleanroom to measure the RoCs and they seem to match the datasheets).

Here's a table comparing the old (design?) RoCs with the new RoCs:

  New RoC Old RoC
LO1 5m 6m
LO2 inf inf
LO3 500mm 750mm
LO4 150mm -450mm
AS1 2m 2.8m
AS2 inf inf
AS3 200mm -2m
AS4 750mm 600mm
PR2 2000m -700m
PR3 1000m 1000m
SR3 1000m -700m

 

The changes looked quite alarming, especially for LO4 and AS3, so I wrote a script to calculate the mode matching between the LO and AS beams called AS_LO_ModeMatching.ipynb and pushed it into the repo. In the notebook a bright AS beam is created by creating a small asymmetry between the arms of ~ 0.003 degrees (~10pm). Amplitude detectors were put at the input ports of the BHD BS to calculate the fields in the AS and LO beams. In particular TEM00, TEM02 and TEM20 were measured for each beam.

The calculation shows that with the old RoCs the mode matching between the LO and AS beams is 99% while for the new RoCs it is ~ 50%.

 

  16765   Thu Apr 7 20:41:15 2022 TommyUpdateElectronicsRFSoC 2x2 Board -- Gain Plotter

In this file (under Tommy), we have a notebook which runs through a spectrum of frequencies and determines the gain response of the attached filter. Below we have the output of a high pass filter. We use IQ demodulation to change IQ componets to DC. Then using a butterworth filter, we read out the DC components and determine the gain's magnitude and phase. However, the phase seems very noisy. This is because the oscillators in the different tiles are independent and a random phase is introduced by changing the mixer frequency in individual tiles. To resolve this we need Multi Tile Synchronization or "MTS". 

Original Pynq Support Forum Query: https://discuss.pynq.io/t/rfsoc-2x2-phase-measurement/3892

We also have the code to fit a resposne function using IIRregular, but this is not as useful without proper phase data.

  16764   Thu Apr 7 20:37:06 2022 TommyUpdateElectronicsRFSoC 2x2 Board -- Simple Tone Generator

In the "Tommy" sub folder, I created a new notebook called "SimpleToneGenerator". This tunes the DAC and ADC mixers to a single frequency and reads off the Time Series and Fourier components. We can alos easily check the demodulation scheme and implement butterworth filters to check their function.

  16763   Thu Apr 7 20:33:42 2022 TommyUpdateElectronicsRFSoC 2x2 board -- setup for remote work

To access the board remotely through the 40m lab ethernet port, use

ssh -N -L localhost:1137:localhost:9090 xilinx@<ip_address>

Then in the browser go to

localhost:1137/lab

Other SSH commands using different ports or without the -N -L seemed to fail to open Jupyter. This way has been successful thereafter.

Quote:

[Tommy, Paco]

Since last week I've worked with tommy on getting the RFSoC 2x2 board to get some TFs from simple minicircuits type filters. The first thing I did was set up the board (which is in the office area) for remote access. I hooked up the TCP/IP port to a wall ethernet socket (LIGO-04) and the caltech network assiggned some IP address to our box. I guess eventually we can put this behind the lab network for internal use only.

After fiddling around with the tone-generators and spectrum analyzer tools in loopback configuration (DAC --> ADC direct connection), we noticed that lower frequency (~ 1 MHz) signals were hardly making it out/back into the board... so we looked at some of the schematics found here and saw that both RF data converters (ADC & DAC) interfaces are AC coupled through a BALUN network in the 10 - 8000 MHz band (see Attachment #1). This is in principle not great news if we want to get this board ready for audio-band DSP.

We decided that while Tommy works on measuring TFs for SHP-200 all the way up to ~ 2 GHz (which is possible with the board as is) I will design and put together an analog modulation/demodulation frontend so we can upconvert all our "slow" signals < 1MHz for fast, wideband DSP. and demodulate them back into the audio band. The BALUN network is pictured in Attachment #2 on the board, I'm afraid it's not very simple to bypass without damaging the PCB or causing some other unwanted effect on the high-speed DSP.

 

  16762   Thu Apr 7 17:59:51 2022 AnchalSummaryBHDPart IIa of BHR upgrade - IR laser alignment on Yarm

[Anchal, Paco, JC, Tega]

Today we have aligned the Yarm cavity for IR, with verified 1.5 roundtrips. We also placed following optics in the BS table.

  • POXM1 (installed earlier, not eloged)
  • SRMOL2
  • POYM1
  • SRMOL1
  • POYM2

We also cleared the in-air BS table of all previous optics. JC and Tega setup HeNe laser for Oplevs roughly for now. Tega also transported the POY RFPD from ITMY in-air table to the BS in-air table. We aligned the POY path to the table, but we had to move ITMY after that to get 2nd roundtrip in the arm cavity, which misaligned our POY path again. POY path would need to be modified tomorrow.

 

  16761   Thu Apr 7 11:47:22 2022 YehonathanUpdateBHDInitial BHD modeling: AS - LO mode matching

I begin modeling the initial BHD setup using Finesse. I started with copying C1_w_BHD.kat from the 40m/bhd repo and making changes to reflect the current BHD setup:

1. OMCs were removed.

2. Only 1 PD per BHD port was left.

3. Transmission of PR2 was changed to 2.2%. The PRG was calculated to be ~15.5.

4. Actual RoCs of new optics were dialed in (Yesterday me and Paco went into the cleanroom to measure the RoCs and they seem to match the datasheets).

Here's a table comparing the old (design?) RoCs with the new RoCs:

  New RoC Old RoC
LO1 5m 6m
LO2 inf inf
LO3 500mm 750mm
LO4 150mm -450mm
AS1 2m 2.8m
AS2 inf inf
AS3 200mm -2m
AS4 750mm 600mm
PR2 2000m -700m
PR3 1000m 1000m
SR3 1000m -700m

 

The changes looked quite alarming, especially for LO4 and AS3, so I wrote a script to calculate the mode matching between the LO and AS beams called AS_LO_ModeMatching.ipynb and pushed it into the repo. In the notebook a bright AS beam is created by creating a small asymmetry between the arms of ~ 0.003 degrees (~10pm). Amplitude detectors were put at the input ports of the BHD BS to calculate the fields in the AS and LO beams. In particular TEM00, TEM02 and TEM20 were measured for each beam.

The calculation shows that with the old RoCs the mode matching between the LO and AS beams is 99% while for the new RoCs it is ~ 50%.

  16760   Wed Apr 6 22:51:47 2022 KojiSummaryBHDPart IIa of BHR upgrade - IR laser alignment on Xarm

[Yuta Koji]

We took out the two coil driver units for PR3 and the incorrect arrangement of the output Rs were corrected. The boxes were returned to the rack.

In order to recover the alignment of the PR3 mirror, C1:SUS_PR3_SUSPOS_INMON / C1:SUS_PR3_SUSPIT_INMON / C1:SUS_PR3_SUSYAW_INMON were monitored. The previous values for them were {31150 / -31000 / -12800}. By moving the alignment sliders, the PIT and YAW values were adjusted to be {-31100 / -12700}. while this change made the POS value to be 52340.

The original gains for PR3 Pos/Pit/Yaw were {1, 0.52, 0.2}. They were supposed to be reduced to  {0.077, 0.04, 0.015}.
I ended up having the gains to be {0.15, 0.1, 0.05}. The side gain was also increased to 50.

----

Overall, the output R configuration for PR2/PR3 are summarized as follows. I'll update the DCC.

PR2 Coil Driver 1 (UL/LL/UR) / S2100616 / PCB S2100520 / R_OUT = (1.2K // 100) for CH1/2/3

PR2 Coil Driver 2 (LR/SD) / S2100617 / PCB S2100519 / R_OUT = (1.2K // 100) for CH3

PR3 Coil Driver 1 (UL/LL/UR) / S2100619 / PCB S2100516 / R_OUT = (1.2K // 100) for CH1/2/3

PR3 Coil Driver 2 (LR/SD) / S2100618 / PCB S2100518 / R_OUT = (1.2K // 100) for CH3

  16759   Wed Apr 6 12:03:51 2022 AnchalSummaryBHDPart IIa of BHR upgrade - IR laser alignment on Yarm

[Anchal, Paco]

After Koji reduced the output resistors on PR2/PR3 coil drivers, we got much better actuation range. We aligned TT1 TT2 again to get beam centered on PR2 and PR3. Then we used only PR2 and PR3 to do the input beam alignment to Y arm cavity. Using access in ETMY chamber, we aligned the input beam parallel to cavity axis. Slight changes were required in ETMY alignment offsets to get first roundtrip in same spot on ITMY. remaining alignment is finer and needs to be done with a help of a reflection photodiode and cameras in the control room. Immediate next step is to setup POY path for locking the Yarm with IR.

Side note: Because of the large PIT correction required in PR3, we found that our upper OSEMs were hitting totally bright limit and lower OSEMs were hitting totally dark limit on PR3. This also destablized our damping loops. We pushed the upper OSEMs slighlty and pulled back the lower OSEMs slightly to get the PD signal in half shadow region again. This worked and our damping loops are stable again. However, we think we should repeat free swing test in future to diagonalize the input matrix for new OSEM positions.

  16758   Wed Apr 6 01:20:48 2022 KojiSummaryBHDPart IIa of BHR upgrade - IR laser alignment on Xarm

PR2/PR3 Output R for fame OSEMS reduced from 1.2K to 1.2K//100Ohm

I put the R=100Ohm for PR3 with the functions of the units mistakenly swapped. This affects imbalanced actuation of PR3 right now as well as too strong SD

PR2 Coil Driver 1 (UL/LL/UR) / S2100616 / PCB S2100520 / R_OUT = (1.2K // 100) for CH1/2/3

PR2 Coil Driver 2 (LR/SD) / S2100617 / PCB S2100519 / R_OUT = (1.2K // 100) for CH3

PR3 Coil Driver 1 (UL/LL/UR) / S2100619 / PCB S2100516 / R_OUT = (1.2K // 100) for CH3 only

PR3 Coil Driver 2 (LR/SD) / S2100618 / PCB S2100518 / R_OUT = (1.2K // 100) for CH1/2/3

----

The output R was reduced from 1.2k to 1.2k//100 = 92 Ohm.

This means that the face coil gains were increased by a factor of 13.

The original gains for PR2 Pos/Pit/Yaw were {0.7, 0.3, 0.2}. To keep the same loop gain, the new gains were supposed to be {0.054, 0.023, 0.015}.
With the new gain, the oscillations were very slowly reduced. Therefore, I increased the gains to have the gain margin of 2. (i.e. increased the gains until I have the oscillation, and then made it half.)
The new values were {0.2, 0.1, 0.05}. The side gain was 20 and unchanged

For PR3 the same operation has been done.

The original gains for PR3 Pos/Pit/Yaw were {1, 0.52, 0.2}. They were supposed to be reduced to  {0.077, 0.04, 0.015}.
The gains were increased to {0.5, 0.1, 0.1}. The side gain was also increased from 1 to 5.

  16757   Tue Apr 5 18:15:06 2022 AnchalSummaryBHDPart IIa of BHR upgrade - IR laser alignment on Xarm

[Paco, Anchal, Ian, JC]

We attempted the alignment of IR beam into the arm cavities. We used PR2 and PR3 (moved manually as well as using cdsutils) and got the YAW aligned pretty good on both X and Y directions. PIT alignment however turned out to be much harder to align. PR2 PR3 didn't have much range, so we zeroed there offset and tried to use TT1, TT2, MMT1, and MMT2 to align the PIT but it would get clipped before reaching BS table if we were to correct for PIT misalignment happening downstream. We concluded that the issue is that one of the PR2, PR3 mirrors have too much PIT offset in equilibrium position. We have requested Koji to change the output resistors in the coil drivers of PR2 and PR3 so that we can correct for the PIT offset in them directly using the coils and reduce load on upstream optics. We have tweaked TT1, TT2, MMT1, and MMT2 positions today, so we do not have the previous reference anymore.

  16756   Mon Apr 4 17:03:47 2022 AnchalSummaryCDSc1susaux2 slow controls acromag chassis fixed and installed

[Anchal, JC, Ian, Paco]

We have now fixed all issues with the PD mons of c1susaux2 chassis. The slow channels are now reading same values as the fast channels and there is no arbitrary offset. The binary channels are all working now except for LO2 UL which keeps showing ENABLE OFF. This was an issue earlier on LO1 UR and it magically disappeared and now is on LO2. I think the optical isolators aren't very robust. But anyways, now our watchdog system is fully functional for all BHD suspended optics.

  16755   Mon Apr 4 15:49:06 2022 TegaUpdateOptical LeversSimplified sketch on BS table

I have updated the BS table using feedback from Koji and Paco and the attached pdf document is the latest iteration.

  16754   Sat Apr 2 15:46:13 2022 ranaSummaryPEMParticle counter setup near BS Chamber

nice - please update the particle counter page in the 40m wiki. Its probably years out of date.

Quote:

I mounted the particle counter over the BS chamber attached to the cable tray as seen in Attachment 1. The signal cable runs through an active 30ft cable to the 1x2 rack. the wire is labeled and runs properly through the cable tray. The particle counter is plugged in at the power strip attached near the cable tray. The power cord is also labeled. 

 

  16753   Fri Apr 1 22:22:29 2022 KojiUpdateOptical LeversSimplified sketch on MC table

Possibility to swap BS and ITMX tables:
BS table, which Tega said MC table, is 2ft x 4ft. The ITMX table is 3ft x 5ft and only the central 2ft x 4ft area is used. The area around the BS table is the narrowest for the east arm. We need at least (2+delta) ft of the hallway width so that we can move the instrument. I'm not yet sure if the ITMX table can be placed there without precise investigation.
 

  16752   Fri Apr 1 17:02:02 2022 KojiUpdateOptical LeversSimplified sketch on MC table

We are supposed to have BS Oplev Beams. We don't like the shallow angle reflections (i.e. AOI>45deg).

The laser is too big but I suspect the other components are too small. So it'd be check the actual size of the components including the optical mounts that are missing on the figure so far.

  16751   Fri Apr 1 14:26:50 2022 TegaUpdateOptical LeversSimplified sketch on MC table

Here is an early sketch of the MC table. 

 

Quote:

I have made an editable draw.io diagram of the planned simplified BHD setup on the ITMY table (see attached).  10 pts = 1 inch.

This is very sketchy but easily adjustable since we are removing everything but the ITMY Oplev from that table.

 

  16750   Fri Apr 1 14:26:19 2022 Ian MacMillanSummaryPEMParticle counter setup near BS Chamber

I mounted the particle counter over the BS chamber attached to the cable tray as seen in Attachment 1. The signal cable runs through an active 30ft cable to the 1x2 rack. the wire is labeled and runs properly through the cable tray. The particle counter is plugged in at the power strip attached near the cable tray. The power cord is also labeled. 

I restarted the particle counter service in the c1psl computer in the /etc/systemd/system/ folder using the commands

sudo systemctl restart particleCounter
sudo systemctl status particleCounter

I cannged the usb hub assigned in the service file to ttyUSB0 which is what we saw the computer had named it.

Checking the channels from this elog show the same particle count as when testing with the buttons and checking the screen. It seems that the channels had been down but are now restarted.

  16749   Thu Mar 31 18:58:16 2022 Ian MacMillanSummaryBHDPart IIa of BHR upgrade - IR laser alignment on Xarm

[Ian, Paco]

We continued to try and lock the IR laser to the x-arm. Last time we adjusted TT2 PR2 and PR3 to adjust the beam to hit the center of the ITMX. Today we continued to adjust the same optics to make the beam hit the center of the beam splitter and make sure the beam is going in the direction of the ITMY. Then removed a mirror that wasn't even bolted down. I put it on the front right corner of the hood with all the other optics (right in front of a suspended mirror). With this unknown mirror removed the path was clear and hit the center of ITMX. From there we went and opened the ETMX chamber and using the IR scope we adjusted ITMX to put the beam on the center of the ETMX. We continued adjusting ETMX and ITMX until we had three reflections. We stopped here because we realized we should set up POX. We think this because any adjustment to set up the POX would mess up the work we had done on the x arm alignment.

  16748   Tue Mar 29 17:35:54 2022 PacoUpdateSUSDamping fix on BS, AS4, PR2, and PR3

[Ian, Paco]

  • We removed the "cheby" filters from AS4, PR2 and PR3 which had been misplaced after copying from the old SUS models. After removing them, the new SOS damped fine. Note that because of the Input matrices, the filters have to be enabled all at once for the MIMO loop to make sense.
  • We also disabled the "Cheby" filter on BS and saw it damp better. We don't understand this yet, but perhaps it's just a consequence of the many changes in the BSC that have rendered this filter obsolete.
  • we also reduced the damping gains on PR2, PR3 and AS4 to prevent overflow values. After the adjustments the optics were damping fine.
  16747   Tue Mar 29 15:07:11 2022 TegaSummaryVACOpening of MC and ETM chambers

[Anchal, Chub, Ian, Paco, Tega]

Today, we opened the MC chamber, ETMX chamber and ETMY chamber.

  16746   Mon Mar 28 16:25:34 2022 Ian MacMillanUpdateSUSETMY SUS Electronics Replacement - Questions

The point of changing the gains was to return the system to its origional state. ie I wanted the over all gain of the physical components to be the same as when we started. From the CDS side of things nothing else should be changed. The damping filters should remain at their origional values. The cts2um filter was changed to counteract a change in the electronics (replacing them). These changes should cancel eachother out. As for the side control, on 3/4/22 koji reduced the output resistors for the 4 face OSEMs but did not change the the SD one. there fore the SD did not need the same adjustment as the others.

Quote:

After Ian updated the cts2um filters for OSEM, shouldn't the damping gains be increased back by factor of 10 to previos values? Was the damping gain for SIDE ever changed? we found it at 250.

Can you explain why gain_offset filter was required and why this wasn't done for the side coil?

Quote:

I updated the gain of the ct2um filter on the OSEMS for ETMY and decreased their gain by a factor of 9 from 0.36 to 0.04.

I added a filter called "gain_offset" to all the coils except for side and added a gain of 0.48.

together these should negate the added gain from the electronics replacement of the ETMY

 

 

  16745   Sat Mar 26 13:04:38 2022 PacoSummaryBHDPart IIa of BHR upgrade - Green laser alignment on Yarm

I came for a bit to check on the "1Y0 burning smell" reported by Koji last evening. I agree that there's a toasty smell around the 1Y0 rack, but it is hard to tell where it comes from exactly. I think we can use the FLIR heat camera, but this will have to wait for Monday when someone with an iphone is around.

I also played a bit with the YAUX alignment (which is still poor) but I didn't try anything different nor did I seem to have more success. This is also an outstanding item for Monday, where maybe the next step is to install the GTRY path for a CCD.

  16744   Fri Mar 25 18:20:28 2022 AnchalSummaryBHDPart IIa of BHR upgrade - Green laser alignment on Yarm

[Anchal, Paco, Shruti]

  • Today we found out that the laser controller at the Yend had it's inputs disconnected for the FAST PZT input and the Slow temperature control input. The cables were also not labeled that were lying nearby.
  • I tracked the cables and connected the Fast input adn the slow input.
  • We then adjusted the temperature slider to go to previously marked laser crystal temperature of 40.1 degrees celsius. But we could not find the the beatnote of Y end IR with teh Main laser IR.
  • We scanned the laser crystal temeprature until we could see a beatntoe with the main IR laser and stopped in that region.
  • Then we continued looking for flashing and HOM, but no luck.
  • It was interesting that when the PDH loop was closed, we saw fluctuations in relfection photodiode DC value, but not so much when the loop was not closed.
  • We tried flipping the sign of the loop with no luck and we tried changing the gain of the loop as well, but no luck.
  • The behavior of this Yend laser and its controller is still suspicious to us. When the loop is closed, the PZT feedback signal should not cause a change in amplitude of the light so much as we see it.
  • Also the fact that when loop is clsoed we do not see signs of flashing in the reflection photodiode is interesting. If anyone has any ideas, that would be good.
  16743   Fri Mar 25 11:39:14 2022 AnchalUpdateSUSETMY SUS Electronics Replacement - Questions

After Ian updated the cts2um filters for OSEM, shouldn't the damping gains be increased back by factor of 10 to previos values? Was the damping gain for SIDE ever changed? we found it at 250.

Can you explain why gain_offset filter was required and why this wasn't done for the side coil?

Quote:

I updated the gain of the ct2um filter on the OSEMS for ETMY and decreased their gain by a factor of 9 from 0.36 to 0.04.

I added a filter called "gain_offset" to all the coils except for side and added a gain of 0.48.

together these should negate the added gain from the electronics replacement of the ETMY

 

  16742   Thu Mar 24 19:20:28 2022 AnchalSummaryBHDPart IIa of BHR upgrade - Green laser alignment on Yarm

[Anchal, Paco]


Uneloged: yesterday, Paco and Ian tried locking the green laser on Yend to the Y-arm. Nothing noteworthy happened, no luck.


  • Started with aligning the green injection to the Iris on table (it was misaligned) and getting it on the ITMY center.
  • Then we realized that the beam is going through the focusing leans very offceter. This makes the beam warped and deflected inside that needs to be corrected by mirrors.
  • So we decided to adjust the lens position first.
  • We realized that lens position is hard to adjust along with keeping the beam hiting ITMY.
  • So, we decided to remove the lens first and just align the green beam directly with the Iris on the table and the ITMY center. The beam became as big as ITMY when it reached there without the lens in path. We aligned it to fall on the mirror.
  • Then we placed the lens back so that the beam goes through the center of the lens.
  • Then we aligned ETMY to get the reflection back to the reflection PD.
  • But then when we checked the beam at ITMY again, it was severely missing ITMY. The possible reason is that ETMY is a diverging lens for the beam as well and changing its position changed the beam direction in the beam tube.
  • We decided to do another iteration of lens positioning with the new ETMY position. So same steps, removed lens, aligned beam with iris and ITMY, placed back the lens, aligned ETMY again to get reflection back on the PD.
  • This time the beam was still reaching ITMY. So we were glad that our iterations converged in 2 repititions.
  • Then we aligned ITMY to get counter propagating beam.
  • At the end of this alignment:
    • We see dips in relfection photodiode counts from 12200 to 9500.
    • The ETMY oplev is partially out of range (only 1 of the quadrant seeing light) but that could be because we changed the lens position. So Oplev servo is off.
  • We confirmed that the PDH servo is properly connected and loop is on. But we see no HOM catching lock for YARM even after us doing about 10 min search with ITMY-ETMY alignments.
  • We are leaving things as it is. We'll continue trying to align the cavity tomorrow. Everyone is welcome to try as long as they restore the positions of the optics to a good state at the end.

 

 

  16741   Mon Mar 21 18:42:06 2022 AnchalSummaryCDSc1auxey1 slow controls issues

Another issue, Green Y Shutter can not be controlled with the EPICS controls right now. This needs to be investigated.

  16740   Mon Mar 21 18:24:07 2022 AnchalSummaryBHDPart IIa of BHR upgrade - Green laser alignment on Yarm

[Anchal, Tega]

We did the following alignment steps on Yarm today:

  • We aligned the green beam going into the ETMY chamber to the three irises on the path.
  • We aligned ETMY to get counter-propagating prompt reflection from ETMY.
  • We were able to see light in the reflection photodiode for the green PDH loop.
  • Then we went to ITMY chamber and checked where the beam is coming to. It was coming very low in height.
  • We walked the beam on the end to get it near the center of ITMY.
  • Then we changed the ETMY alignment again to get back the reflection beam on the reflection photodiode.
  • Then we went again to ITMY chamber and changed the ITMY alignment to get counter-propagating reflection from ITMY.

After this, we came back to the control room. One peculiar thing to note is that the C1:LS-Y_REFL_DC_OUT channel is inverted i.e. it is showing negative values for the reflection DC voltage. On this signal, we see a little bit higher-order mode flashing but it is not bright enough to be seen on the face camera of the suspension. We'll continue aligning the cavity using CDS now to get TEM00 mode.


After trying for a bit, we were able to get flash of about 2000 counts which is about 16% of the max value of 12200 counts. We adjusted the ITMY angular position using the IFO_ALIGN screen but used the OPTICAL_ALIGN offset screen to adjust the angular position of ETMY. The pitch and yaw values for ITMY are 1.0360 and -0.260 respectively whereas the pitch and yaw values for ETMY are 5664.0 and -4937.0 respectively. 

  16739   Mon Mar 21 18:00:05 2022 Ian MacMillanUpdateBHDXARM AUX alignment cont.

[Ian, Paco]

Continuing with the previous alignment that we stoped on friday, we re set up my heavily cleaned iPhone on FaceTime. The Phone alowed us to see the laser on the ITMX and center it on that optic.

  • We increased the trigger level for the watchdog on the optics.
  • Green beam was centered on th ITMX.
  • aligned the green laser: prompt reflection off of the ETMX to maximize the signal on the reflection PD
  • Went back to ITMX chamber, looked for the beam in the tube to see if the laser was hitting the beam tube off of the ITMX. From there we comunicated how to adjust the beam to move the beam towards the ETMX.
  • On the ETMX the reflection PD signal was measured and the ITMX was adjusted using CDS until ripples/"flashes" were seen in the reflection PD channel (C1:ALS-X_REFL_DC_OUT_DQ).
  • Once we saw flashing we tried to minimize the reflection pd signal using only ITMX alignment slider so that all the light was being held in the cavity.
  • Once we had this going the PDH lock engaged the higher order mode that we were seeing.
  • We removed the phone setup and closed up the ITMX chamber.
  • From here we moved to the control room and continude to adjust the ITMX and ETMX to see if we could lock to a lower order mode.
    • This proved unsuccessful so we will resume by trying to replicate the ASS action manually (with M1 - M2 PZT sliders and the REFL PD)
  16738   Mon Mar 21 14:22:52 2022 AnchalUpdateSUSETMY Alignmnet offsets needed to be changed

I'm not sure why but the PIT and YAW offset values of +2725 and -2341 were not sufficient for the reflected OPLEV beam to reach the OPLEV QPD. I had to change the C1:SUS-ETMY_PIT_OFFSET to 5641 and C1:SUS-ETMY_YAW_OFFSET to -4820 to come back to center of the OPLEV QPD. We aligned the Oplevs to center before venting, so hopefully this is our desired ETMY position.

On another note, the issue of ETMY unable to damp was simple. The alignment offsets were on to begin with with values above 1000. This meant that whenever we enabled coil output, ETMY would necessarily get a kick. All I had to do was keep alignment offsets off before starting the damping and slowly increase the alignment offsets to desired position.

Quote:
 

- This modification allowed me to align the oplev spot to the center of the QPD. C1:SUS-ETMY_PIT_OFFSET and C1:SUS-ETMY_YAW_OFFSE are +2725 (8%FS) and -2341 (7%FS), respectively.

  16737   Fri Mar 18 19:10:51 2022 AnchalSummaryCDSc1auxey1 slow controls issues

I started the modbusIOC service on c1auxey1 and added PD variance channels for UR and SD as well.  There are unfortunately two issues here:

  • The enable monitors are reading NOT of what they should read. The optical isolator circuit might need to be changed.
  • ETMY is not damping now. This is strange and was seen in the use to other acromag chassis as well where AS4 and PR2 are unable to damp. This is weird since the acromag chassis are not part of the damping loop, maybe it is a coincidence. Next time we should check if we still have this issue when acromag chassis is disconnected from ETMY.

 

  16736   Fri Mar 18 18:39:13 2022 YehonathanSummaryCDSc1auxey1 slow controls acromag chassis installed, powered

{Yehonathan, Anchal}

We connected the c1auxey1 chassie to the different boxes (coil drivers, SAT amp, etc.) using DB9 cables and labeling them in the process. We ran out of 2.5 foot DB9 cables so we used 5 foot as a temporary solution.

The chassie was powered, but a two issues arised:

1. The Acromags didn't turn on.

2. When connecting the green laser shutter BNC cable, the power supply overloaded.

We took the chassie back to the bench. The wire that powers the Acromags was disconnected. We made a new longer wire and made sure it is not connected flimsily.

The issue with the BNC turned out to be a much deeper problem: The GND and EXC wires on the DIN rail connector were switched! Making the shield of the BNC to have high volatage compared to the shield of the green shutter causing current to overflow when the BNC was connected.

We switched back the EXC and GND wires. Not trusting the digital I/O tests that were done before due to this mistake we tested some of the I/Os using a spare coil driver. We tested both the inputs and the outputs and they all seemed to work.

Finally, we also noticed that the 2 RTS DB9s were wrongly female type so we switched them to males. We closed the lead on chassie and installed it back in the rack. We connected the cables and saw that the green shutter BNC cable was no longer shorting the power supply.

  16735   Fri Mar 18 17:15:19 2022 PacoUpdateBHDXARM AUX alignment cont.

[Paco, Ian]

First, we re-balanced the ITMX OSEMs so that they would damp at around half-a-shadow.


Then, we set up a clean camera inside ITMX chamber looking at the ITMX optic. Then, using the live feed we aligned the AUX beam from the ETMX station using M1 and M2. The camera was great to help us align the beam properly close to the ITMX center. It wasn't very long until we could see a green beam on the IR card, but we didn't really see any flashing, so this may just be the bare transmission away from XARM resonance (Attachment #1).

Ian checked the reflection from ITMX using the IR card with holes, and he pretty much only saw one beam spot, so we turned to look for a beam scattering on the vacuum tube but didn't really see anything. This could mean that we were hitting the ETMX again, or missing slightly, or missing completely. We tried scanning the ITMX pitch and yaw using the bias (alignment) sliders, and with the illuminators off, try seeing some scattered green beam on the ETMX. We can't really see anything yet, but we will keep trying. If there are any tips on our method, it would be great to know them.

  16734   Thu Mar 17 19:12:44 2022 AnchalSummaryCDSc1auxey1 slow controls acromag chassis installed, not powered

[Anchal, Tega]

We installed c1auxey1 computer and the acromag chassis in 1Y4. The computer has been configured properly for nfs mounts to happen and we have initialized a git repo for /cvs/cds/caltech/target/c1auxey1 directory which stores all files for running modbusIOC service on this computer. We connected 18V power source but have not connected the 24V power yet  as we need to make a new connector for it. Going on what Koji recommended, we'll connect the 24V power input to 18 V strip as well as the acromags can run on that voltage too.

  16733   Thu Mar 17 17:23:22 2022 AnchalSummaryBHDPart IIa of BHR upgrade - Green laser alignment

[Anchal, Yehonathan] (on Y Arm)

We first checked if the PZT mirrors M1 and M2 can be controlled. They indeed show no motion even after being connected with a power supply. So green injection path can not be aligned using cds controls right now.

We also noticed that all ETMY slow controls and monitors are offline. That's because the electronics upgrade did not include acromag chassis. This means that the DC bias adjustment is not accessible for ETMY.

Alignment work:

  • We first aligned the green injection beam to cross two irises in the injection path to ETMY chamber.
  • Then I (Anchal) went inside the ITMY chamber to find the green light. Yehonathan controlled the injection path while I gave feedback from ITMY side. We aligned injection path to get the beam near center of ITMY.
  • Then I aligned ITMY through sitemap>IFO>Align to get counter propagating reflection near the ITMY side.
  • We were able to see reflection from ETMY hitting the beam tube.
  • Since DC alignment controls of ETMY are not accessible, we used the Alignment offset in rtcds model which puts dc offset in the coil driver to get the reflected beam from ETMY to come to ITMY table, about 1 inch above the table and about 3 inch south of ITMY SOS.
  • But we got limitted by the DAC overflow on ETMY at this point. Several back and forth attempts to relief ETMY were unsuccesful.

Possible issues:

  • We think having the HV coil driver set up for ETMY is important for this alignment work if not essential. The coil drivers of ETMY are near saturation.
  • We also think that addition of two new suspensions in ITMY table and then counter weights to balance the table, has depressed table height slightly. We need to work out how to change injection path height and angle accordingly.

[Paco, Ian] (on X Arm)

http://nodus.ligo.caltech.edu:8080/40m/16732

  16732   Thu Mar 17 16:50:53 2022 PacoUpdateBHDXARM AUX alignment

[Ian, Paco]

We opened the ITMX chamber and

  • Inspected PR2 and LO1 and didn't see anything suspicious.
  • Took the EQ stops off ITMX in preparation for alignment

Alignment:

  • Ian started at the ETMX station, slightly tweaking the M1 M2 XAUX mirrors by hand and Paco on the ITMX chamber looking at the beam. The beam was visibly making it to the ITMX optic and reflecting back though at a negative pitch angle. No small motion on the M1-M2 knobs caused visible correction on this.
  • We decided to try adding an alignment offset to ITMX, and quickly saturated to 25000 (on the SUS screen) but fell short of fully correcting this PIT offset.
  • We scanned further the M1 M2 input alignment, and briefly lost the first reflection.
    • We recovered the first reflection, but the beam spot is now incident near the LL OSEM position.
  • We kept scanning M1 - M2 with Ian on the ITMX chamber this time, providing feedback through facetime.
  • Leaning on the ITMX chamber table, we noticed the magnitude of PIT correction left to be done and verified that M1-M2 two axis alignment + ITMX offsets should be enough for it.

We stopped our effort here, the XAUX beam spot is near the lower half of the ITMX face. Tomorrow, we will resume, but we will use airpods and a clean go-pro for real-time audiovisual feedback. Furthermore, ITMX OSEMs should be rebalanced as they haven't been touched after the table was balanced for PR2 and LO1.

  16731   Thu Mar 17 11:40:41 2022 PacoUpdateSUSETMY green PZT HV supply

[Anchal, Paco]

We installed a HV kepco power supply for the green PZT steering the YAUX beam. We did this in anticipation of the YARM alignment to take place this afternoon. We borrowed an unused DC power supply labeled "OMC Power spply" and made an appropriate SMA connection (Attachment #1), Then we set the Vset to 100.0 Volts and the current limit to 10 mA. Once we enabled the output we saw the 5.6 mA of current drawn by the eurocard in accordance with the wiki log (Attachment #2). 

It may not be possible to use the PZT as per this so this work may not have a direct impact on our upcoming alignment task.


We probably bumped the ethernet cable (martian network) on c1iscey, so the FE models stopped showing up on the medm screens. When we connected it back, it seemed like the FE kept running the model and only IPC showed error. We restarted the rtcds models on c1iscey and burtrestored to today morning 5:19 am.

burtwb -f /cvs/cds/rtcds/caltech/c1/burt/autoburt/today/05:19/c1scyepics.snap -l /tmp/controls_1220317_115006_0.write.log -o /tmp/controls_1220317_115006_0.nowrite.snap -v

ETMY is properly damping now and the oplev is roughly centered as well but the OPLEV Servos are off. We did not turn them on.
We should be able to carry out our cavity arm alignment today afternoon on both arms now.

  16730   Tue Mar 15 18:45:12 2022 AnchalSummaryBHDPart X of BHR upgrade - BHDBS Path setup

[Paco, Anchal]

BS Chamber work

  • ASL was positioned in nominal place.
  • PR3 was moved to its nominal place from temprorary position.
  • BS Table was rebalanced
  • Earthquake stops were removed from all SOS from BS table (LO2, SR2, BS, PR3)

ITMY Chamber work

  • AS2, AS3, LO3, LO4, and BHDBS were positioned in the nominal place.
  • AS1 was moved to its nominal place from temporary position.
  • ITMY tbale was rebalanced
  • Earthquake stops were removed from all SOS from ITMY table (AS1, AS4, ITMY)
  16729   Tue Mar 15 18:42:37 2022 AnchalSummaryBHDPart IX of BHR upgrade - AS1 resuspended and OSEMs tuned

http://nodus.ligo.caltech.edu:8080/40m/16722

http://nodus.ligo.caltech.edu:8080/40m/16716

  16728   Tue Mar 15 14:10:41 2022 AnchalSummaryCDSc1su2 model remade, reinstalled, restarted after the update

I have restarted c1su2 model with the connections of Run Acquire switch to analog filters on coil drivers. Following steps were taken:

First ssh to c1sus2 and then:

controls@c1sus2:~ 0$ rtcds make c1su2
buildd: /opt/rtcds/caltech/c1/rtbuild/release
### building c1su2...
Cleaning c1su2...
Done
Parsing the model c1su2...
Done
Building EPICS sequencers...
Done
Building front-end Linux kernel module c1su2...
Done
RCG source code directory:
/opt/rtcds/rtscore/branches/branch-3.4
The following files were used for this build:
/opt/rtcds/userapps/release/cds/common/models/lockin.mdl
/opt/rtcds/userapps/release/cds/common/models/rtbitget.mdl
/opt/rtcds/userapps/release/cds/common/models/rtdemod.mdl
/opt/rtcds/userapps/release/isc/common/models/QPD.mdl
/opt/rtcds/userapps/release/sus/c1/models/c1su2.mdl
/opt/rtcds/userapps/release/sus/c1/models/lib/sus_single_control.mdl

Successfully compiled c1su2
***********************************************
Compile Warnings, found in c1su2_warnings.log:
***********************************************
WARNING  *********** No connection to subsystem output named  SUS_DAC1_12  
WARNING  *********** No connection to subsystem output named  SUS_DAC1_13  
WARNING  *********** No connection to subsystem output named  SUS_DAC1_14  
WARNING  *********** No connection to subsystem output named  SUS_DAC1_15  
WARNING  *********** No connection to subsystem output named  SUS_DAC2_7  
WARNING  *********** No connection to subsystem output named  SUS_DAC2_8  
WARNING  *********** No connection to subsystem output named  SUS_DAC2_9  
WARNING  *********** No connection to subsystem output named  SUS_DAC2_10  
WARNING  *********** No connection to subsystem output named  SUS_DAC2_11  
WARNING  *********** No connection to subsystem output named  SUS_DAC2_12  
WARNING  *********** No connection to subsystem output named  SUS_DAC2_13  
WARNING  *********** No connection to subsystem output named  SUS_DAC2_14  
WARNING  *********** No connection to subsystem output named  SUS_DAC2_15  
***********************************************
controls@c1sus2:~ 0$ rtcds install c1su2
buildd: /opt/rtcds/caltech/c1/rtbuild/release
### installing c1su2...
Installing system=c1su2 site=caltech ifo=C1,c1
Installing /opt/rtcds/caltech/c1/chans/C1SU2.txt
Installing /opt/rtcds/caltech/c1/target/c1su2/c1su2epics
Installing /opt/rtcds/caltech/c1/target/c1su2
Installing start and stop scripts
/opt/rtcds/caltech/c1/scripts/killc1su2
/opt/rtcds/caltech/c1/scripts/startc1su2
Performing install-daq
Updating testpoint.par config file
/opt/rtcds/caltech/c1/target/gds/param/testpoint.par
/opt/rtcds/rtscore/branches/branch-3.4/src/epics/util/updateTestpointPar.pl -par_file=/opt/rtcds/caltech/c1/target/gds/param/archive/testpoint_220315_135808.par -gds_node=26 -site_letter=C -system=c1su2 -host=c1sus2
Installing GDS node 26 configuration file
/opt/rtcds/caltech/c1/target/gds/param/tpchn_c1su2.par
Installing auto-generated DAQ configuration file
/opt/rtcds/caltech/c1/chans/daq/C1SU2.ini
Installing Epics MEDM screens
Running post-build script

/opt/rtcds/userapps/release/cds/common/scripts/generate_KisselButton.py 4 5 C1:SUS-AS1_INMATRIX > /opt/rtcds/caltech/c1/medm/c1su2/C1SUS_AS1_INMATRIX_KB.adl
/opt/rtcds/userapps/release/cds/common/scripts/generate_KisselButton.py 2 4 C1:SUS-AS1_LOCKIN_INMTRX > /opt/rtcds/caltech/c1/medm/c1su2/C1SUS_AS1_LOCKIN_INMTRX_KB.adl
/opt/rtcds/userapps/release/cds/common/scripts/generate_KisselButton.py 5 6 C1:SUS-AS1_TO_COIL --fi > /opt/rtcds/caltech/c1/medm/c1su2/C1SUS_AS1_TO_COIL_KB.adl
/opt/rtcds/userapps/release/cds/common/scripts/generate_KisselButton.py 4 5 C1:SUS-AS4_INMATRIX > /opt/rtcds/caltech/c1/medm/c1su2/C1SUS_AS4_INMATRIX_KB.adl
/opt/rtcds/userapps/release/cds/common/scripts/generate_KisselButton.py 2 4 C1:SUS-AS4_LOCKIN_INMTRX > /opt/rtcds/caltech/c1/medm/c1su2/C1SUS_AS4_LOCKIN_INMTRX_KB.adl
/opt/rtcds/userapps/release/cds/common/scripts/generate_KisselButton.py 5 6 C1:SUS-AS4_TO_COIL --fi > /opt/rtcds/caltech/c1/medm/c1su2/C1SUS_AS4_TO_COIL_KB.adl
/opt/rtcds/userapps/release/cds/common/scripts/generate_KisselButton.py 4 5 C1:SUS-LO1_INMATRIX > /opt/rtcds/caltech/c1/medm/c1su2/C1SUS_LO1_INMATRIX_KB.adl
/opt/rtcds/userapps/release/cds/common/scripts/generate_KisselButton.py 2 4 C1:SUS-LO1_LOCKIN_INMTRX > /opt/rtcds/caltech/c1/medm/c1su2/C1SUS_LO1_LOCKIN_INMTRX_KB.adl
/opt/rtcds/userapps/release/cds/common/scripts/generate_KisselButton.py 5 6 C1:SUS-LO1_TO_COIL --fi > /opt/rtcds/caltech/c1/medm/c1su2/C1SUS_LO1_TO_COIL_KB.adl
/opt/rtcds/userapps/release/cds/common/scripts/generate_KisselButton.py 4 5 C1:SUS-LO2_INMATRIX > /opt/rtcds/caltech/c1/medm/c1su2/C1SUS_LO2_INMATRIX_KB.adl
/opt/rtcds/userapps/release/cds/common/scripts/generate_KisselButton.py 2 4 C1:SUS-LO2_LOCKIN_INMTRX > /opt/rtcds/caltech/c1/medm/c1su2/C1SUS_LO2_LOCKIN_INMTRX_KB.adl
/opt/rtcds/userapps/release/cds/common/scripts/generate_KisselButton.py 5 6 C1:SUS-LO2_TO_COIL --fi > /opt/rtcds/caltech/c1/medm/c1su2/C1SUS_LO2_TO_COIL_KB.adl
/opt/rtcds/userapps/release/cds/common/scripts/generate_KisselButton.py 4 5 C1:SUS-PR2_INMATRIX > /opt/rtcds/caltech/c1/medm/c1su2/C1SUS_PR2_INMATRIX_KB.adl
/opt/rtcds/userapps/release/cds/common/scripts/generate_KisselButton.py 2 4 C1:SUS-PR2_LOCKIN_INMTRX > /opt/rtcds/caltech/c1/medm/c1su2/C1SUS_PR2_LOCKIN_INMTRX_KB.adl
/opt/rtcds/userapps/release/cds/common/scripts/generate_KisselButton.py 5 6 C1:SUS-PR2_TO_COIL --fi > /opt/rtcds/caltech/c1/medm/c1su2/C1SUS_PR2_TO_COIL_KB.adl
/opt/rtcds/userapps/release/cds/common/scripts/generate_KisselButton.py 4 5 C1:SUS-PR3_INMATRIX > /opt/rtcds/caltech/c1/medm/c1su2/C1SUS_PR3_INMATRIX_KB.adl
/opt/rtcds/userapps/release/cds/common/scripts/generate_KisselButton.py 2 4 C1:SUS-PR3_LOCKIN_INMTRX > /opt/rtcds/caltech/c1/medm/c1su2/C1SUS_PR3_LOCKIN_INMTRX_KB.adl
/opt/rtcds/userapps/release/cds/common/scripts/generate_KisselButton.py 5 6 C1:SUS-PR3_TO_COIL --fi > /opt/rtcds/caltech/c1/medm/c1su2/C1SUS_PR3_TO_COIL_KB.adl
/opt/rtcds/userapps/release/cds/common/scripts/generate_KisselButton.py 4 5 C1:SUS-SR2_INMATRIX > /opt/rtcds/caltech/c1/medm/c1su2/C1SUS_SR2_INMATRIX_KB.adl
/opt/rtcds/userapps/release/cds/common/scripts/generate_KisselButton.py 2 4 C1:SUS-SR2_LOCKIN_INMTRX > /opt/rtcds/caltech/c1/medm/c1su2/C1SUS_SR2_LOCKIN_INMTRX_KB.adl
/opt/rtcds/userapps/release/cds/common/scripts/generate_KisselButton.py 5 6 C1:SUS-SR2_TO_COIL --fi > /opt/rtcds/caltech/c1/medm/c1su2/C1SUS_SR2_TO_COIL_KB.adl
safe.snap exists
controls@c1sus2:~ 0$

Then on rossa, run activateSUS2DQ.py which creates a file C1SU2.ini.NEW. Remove old backup file C1SU2.ini.bak, rename C1SU2.ini to C1SU2.ini.bak and rename C1SU2.ini.NEW to C1SU2.ini:

~> cd /opt/rtcds/caltech/c1/chans/daq/
daq>python2 activateSUS2DQ.py 
/opt/rtcds/caltech/c1/chans/daq/C1SU2.ini
daq>rm C1SU2.ini.bak
daq>mv C1SU2.ini C1SU2.ini.bak
daq>mv C1SU2.ini.NEW C1SU2.ini

Then ssh back to c1sus2 and restart the rtcds model:

controls@c1sus2:~ 0$ rtcds restart c1su2
### stopping c1su2...
### starting c1su2...
c1su2epics: no process found
Number of ADC cards on bus = 2
Number of DAC16 cards on bus = 3
Number of DAC18 cards on bus = 0
Number of DAC20 cards on bus = 0
Specified filename iocC1.log does not exist.
c1su2epics C1 IOC Server started
c1su2 RT ready in 4
awg_server Version $Id$
channel_client Version $Id$
testpoint_server Version $Id$
/opt/rtcds/caltech/c1/target/gds/bin/awgtpman -s c1su2 -l /opt/rtcds/caltech/c1/target/gds/awgtpman_logs/c1su2.log started on host c1sus2 hostid ffffffffa8c05771 
awgtpman Version $Id$
controls@c1sus2:~ 0$

Then restart daqd services from rossa and burtrestore to latest snap of c1su2epics.snap:

daq>telnet fb 8083
Trying 192.168.113.201...
Connected to fb.martian.
Escape character is '^]'.
daqd> shutdown
OK
Connection closed by foreign host.
daq>burtgooey
>burtwb -f /opt/rtcds/caltech/c1/burt/autoburt/latest/c1su2epics.snap -l /tmp/controls_1220315_140755_0.write.log -o /tmp/controls_1220315_140755_0.nowrite.snap -v <
daq>

All suspensions are back online and everything is same as before now. Will test later the Run/Acquire switch functionality.

  16727   Tue Mar 15 13:49:44 2022 Ian MacMillanBureaucracyTreasureNew Screwdriver Bits

I have received the new screwdriver bits that will work with the two electric screwdrivers we have. I have distributed them in the 40m. Some are in the electronics stations and some are in the toolbox in the lab. The new electric screwdriver (which looks like a drill but takes typical screwdriver bits) is in the room with the workshop. It is in the blue Makita box. For some reason, lots of the old bits were rounded because of incorrect use. I have thrown the unuseable ones out.

I also requested some screw extractors in case we need them. the one we have now is really big and may not work on smaller screws.

  16726   Tue Mar 15 11:52:34 2022 AnchalSummaryCDSc1su2 model updated for sending Run/Acquire Binary Output to Binary Interface card

I routed the XXX_COIL_DW signals from the 7 SOS blocks in c1su2.mdl (located at /cvs/cds/rtcds/userapps/trunk/sus/c1/models/c1su2.mdl) to the binary outputs from the FE model. The routing is done such that when these binary outputs are routed through the binary interface card mounted on 1Y0, they go to the acromag chassis just installed and from there they go to the binary inputs of the coil drivers together with the acromag controlled coil outputs.

I have not restarted the rtcds models yet. This needs more care and need to follow instructions from 40m/16533. Will do that sometime later or Koji can follow up this work.

  16725   Tue Mar 15 10:45:31 2022 PacoUpdateGeneralAssembled small in-vac optics

[Paco]

This morning I assembled LO3, LO4 and AS3 (all mirrors) onto polaris K1 mounts. The mounts stand as per this elog, on 4.5" posts with 0.5" Al spacers to match the beam heigth of 5.5". I also assembled ASL by adding a 0.14" Al spacer, and finally, recycled two DLC mounts (from the XEND flowbench) and posts to mount the 2 inch diameter beamsplitters BHDBS and AS2 (T=10%). I stored the previous 2" optics in the CVI and lambda optic cases and labeled appropriately.

  16724   Mon Mar 14 12:20:05 2022 AnchalSummaryCDSc1susaux2 slow controls acromag chassis installed

[Anchal, Yehonathan, Ian]

We installed c1susaux2 acromag chassis in 1Y0 with c1susaux2 computer. We connected PD monitors, Binary inputs, Binary outputs, and Run/Acquire RTS signals for 6 of the 7 suspensions. We ran out of DB9 cables to connect PR3. Of the ones that were connected, LO2, AS1, AS4, SR2, and PR2 are showing no issues in the functionality from the chassis. For LO1, everything is working except for UR EnableMon channel. The enable monitor does not show an ON state for the coil even though the coil driver chassis shows that it is ON via the LED lights. A possible reason could be that a wire got disconnected when we closed the chassis (there are a lot of wires pushing against each other. Another reason could be that the optical isolator ISO10 could have developed a bad channel on channel 2. The circuit was tested before closing the chassis, so not sure what went wrong after closing it.

PR2 is showing a non-acromag chassis related issue. As soon as we close the loop by enabling the coils, the watchdog triggers because the loop is unstable. Not sure what has changed for PR2, but someone should take a look at it.

For the issue with LO1, I suggest we keep a note that the C1:SUS-LO1_UR_ENABLEMon channel is faulty and don't take its value seriously. We should diagnose and fix this issue once we have more reasons to disconnect the chassis and open it.

 

  16723   Fri Mar 11 16:43:03 2022 Ian MacMillanUpdateSUSETMY SUS Electronics Replacement

I updated the gain of the ct2um filter on the OSEMS for ETMY and decreased their gain by a factor of 9 from 0.36 to 0.04.

I added a filter called "gain_offset" to all the coils except for side and added a gain of 0.48.

together these should negate the added gain from the electronics replacement of the ETMY

  16722   Thu Mar 10 10:05:49 2022 PacoUpdateSUSAS1 free swing test

[Paco, Ian]

  • Begin free swinging test for AS1 at 10:05 AM, set for ~ 2.04 hours.
    • Test failed because damping failed to disable.
  • Restart free swinging test for AS1 at 15:06, set for 2.04 hours.
    • Success (Attachment #1 shows the DOF input matrix diagonalization effect)

Of slight concern is the side to other degrees of freedom coupling, but this is definitely an improvement from last time.

  16721   Thu Mar 10 09:39:59 2022 JordanUpdateVACLeak Testing of New Manual Gate Valve - Attempt #2

This morning Chub and I leak checked the manual gate valve with the RGA and helium. There was no change in the helium partial pressure while spraying helium around the flanges, all looks good.

 

I also took a 100 AMU analog scan, after the filament had warmed up overnight and the plot was quite noisy even with the lowest scan speed. I recommend this unit go back to SRS for a filament replacement/recalibration. I am worried yesterday's "vent" of the RGA volume may have burned the filament. See the comparison of yesterday's analog scan to today's below.

  16719   Wed Mar 9 12:57:52 2022 KojiUpdateBHDSimplified BHD readout sketch on ITMY table

- BHD beams were already mixed in the chamber. So we don't need a BS on the table. (Probably there is no BS already)

- We don't need to split each BHD beam. One PD per BHD beam is OK for now.

- Check if the BHD paths have reasonable angles from the windows so that the beams do not hit the chamber wall.

- We need the POY path. POY indeed goes to the BS table

  16718   Wed Mar 9 11:52:18 2022 YehonathanUpdateBHDSimplified BHD readout sketch on ITMY table

I have made an editable draw.io diagram of the planned simplified BHD setup on the ITMY table (see attached).  10 pts = 1 inch.

This is very sketchy but easily adjustable since we are removing everything but the ITMY Oplev from that table.

  16717   Wed Mar 9 10:23:28 2022 JordanUpdateVACLeak Testing of New Manual Gate Valve - Attempt #1

Jordan, Chub, Paco

Chub and I went into the lab this morning to leak check the new gate valve after pumping over the weekend. This would be done through the RGA and spraying helium around the newly installed flanges. The RGA is set to monitor the partial pressure of helium versus time and we visually watch for any spikes in pressure to indicate an air leak.

So, I mistakenly thought the gate valve was opened and both sides were being pumped on, this was not the case. The vavle was closed so there was a pressure differential whenI turned the handle it tripped the interlocks and closed V4. I closed VM3 to the RGA volume to prevent the filament from being damaged.

Then the medm screen for the vac controls started flashing rapidly, I closed the window and reopened the controls to find all the panels were white, but we could still see the read only vac screen. So Paco restarted c1vac and the controls were restored. V7 closed, but was then reopened.

We now need to restart pumping with the odd pressure differentials between V4 and VM3.

Plan

- TP2 turned off along with the AUX pump

- VM3 opened to vent RGA volume (RGA turned off, filament cooled)

- Open the manual vent screw on TP1 to bring the RGA+TP1 volume back to atmosphere, now no pressure diff. on V4

- Open V4, restart AUX pump to rough out the volume

- Connect RP1/3 line to the pump scroll, turn on RP1/3 RGA+TP1 volume went to mtorr

- Slowly open the manual gate valve

- Connect AUX pump to TP2 and rough out

- Restart TP2, once at speed open V4, close V6 and turn off RP1/3

Now we are pumping on the RGA/TP1 volume with TP2, leak check attempt #2 will happen tomorrow morning

ELOG V3.1.3-